OpenStudy (anonymous):

What is (1.2 × 10^–2) × (3 × 10^5)? 3.6 × 10^–4 3.6 × 10^–3 3.6 × 10^3 3.6 × 10^4 What is 8.25 × 10^6 written in standard form 82,500 825,000 8,250,000 82,500,000

OpenStudy (anonymous):

2 question is a than first is i want to say c

OpenStudy (anonymous):

On the first question?

OpenStudy (pixiedust1):

Well, 1.2*3=3.6 10^-2*10^5 -2+5=3 So, we are now left with 3.6*10^3

OpenStudy (pixiedust1):

So yes, C would be correct

OpenStudy (anonymous):

Ok thanks so much!! What about 2nd one?

OpenStudy (pixiedust1):

And for #2, 8.25*10^6=8,250,000

OpenStudy (pixiedust1):

Standard form just means regular, like written out. Nothing fancy.

OpenStudy (anonymous):

Ok thanks you're such a big help!!!!!!! ☺

OpenStudy (pixiedust1):

Not a problem! Need help on any more?

OpenStudy (anonymous):

Nope I am good!

OpenStudy (pixiedust1):

Great! Have an awesome day! :D

OpenStudy (anonymous):

You too!